OpenStudy (anonymous):

Prove the following limit statement: lim (X^2-1)/(x-1) = 2 x-->1

OpenStudy (anonymous):

By using the precise definition of a limit.

OpenStudy (anonymous):

then let's let \( \varepsilon > 0 \) and be done with it ^.^

OpenStudy (anonymous):

sorry @OneMathCat please carry on ^_^

OpenStudy (anonymous):

First of all, note that (x^2-1)/(x-1) = (x+1)(x-1)/(x-1) = x+1, providing x is not equal to 1. Thus, as x approaches 1, x+1 --> 2. So, surely sounds like it's true! Now, for the precise definition, you must show that for all epsilon > 0, there exists delta > 0, such that if 0 < |x-1| < delta, then | (x+1) - 2 | < epsilon. Note that | (x+1) - 2 | = | x - 1| . I'll let you take it from here! Good luck! (So think --- how should you choose delta?)